OpenStudy (sbuck98):

Someone help me answer these 4 problems, I will award a medal, and coins.

OpenStudy (jimmykim):

What r the questions

OpenStudy (sbuck98):

1.) Slope (m)= change in y over change in x. M= y2-41 over x2-x1 show your work on how you found the slope.

OpenStudy (sbuck98):

2. Find the equation using the point slope formula. Point slope formula is y-y1=m(x-x1)

OpenStudy (sbuck98):

3. Give a example of a line parallel to your answer from question #2.

OpenStudy (sbuck98):

4. Give a example of a line perpendicular to your answer from #2.